Fawns between 1 and 5 months old have a body weight that is approximately normally distributed with mean μ = 27.7 kilograms and standard deviation σ = 4.4 kilograms. Let x be the weight of a fawn in kilograms.
A) x < 30 then z < __ is what?
b) 19 < x then __ < z is what?

Found 2 solutions by stanbon, Boreal:
Answer by stanbon(75887)

Fawns between 1 and 5 months old have a body weight that is approximately normally distributed with mean μ = 27.7 kilograms and standard deviation σ = 4.4 kilograms. Let x be the weight of a fawn in kilograms.
A) x < 30 then z < __ is what?
z(30) = (30-27.7)/4.4 = 0.5227
Ans:: z < 0.5227
------------------
b) 19 < x then __ < z is what?
z(19) = (19-27.7)/4.4 = -1.9773
Ans: z > -1.9773
---------------
Cheers,
Stan H.

Answer by Boreal(15235)

z < (x-mean)/sd
=(30-27.7)/4.4<0.68; probability is 0.7523
================
19 z>(19-27.7)/4.4>-1.98
-1.98 < z
probability is 0.9762; the probability it weighs more than 19 kg is 0.9762.